#25 This question involves to find parallel and perpendicular lines.
Find the equation of the line containing the point (4,1) and perpendicular to the line y=-3x+4.
I believe I do y-y1=m(x-x1)
y-1=4(x-4)
y=4x-16+1
y=x-17/4
However, I know this is wrong. the answer is supposed to be y=1/3x-1/3. I have no idea what I need to do. Any help would be appreciated.

Answer by Alan3354(69443)

Find the equation of the line containing the point (4,1) and perpendicular to the line y=-3x+4.
----------
The slope of the line is -3
The slope of lines perpendicular is the negative inverse, +1/3
------------
Use y = mx+b and the point to find b:
1 = (1/3)*4 + b
b = -1/3
y = (1/3)x - 1/3